Find the number of real solutions using the discriminant
x^2 - 25 = 0
I already tried doing this and got two real solutions, but it was marked wrong. Please show your working so I can see where I went wrong. ^_^

Answer by Edwin McCravy(20054)

x^2 - 25 = 0

There is no term in x, so we put 0x in as a place holder.

x^2 + 0x - 25 = 0

So a=1, b=0, and c=-25

Discriminant = b^2-4ac = 0^2-4(1)(-25) = 0+100 = 100

100 is positive so there are 2 real solutions.

Edwin